# CMB delensing with deep learning

### v2

@@ 2.2 The CMB temperature and lensing potential
To acquire additional datasets and account for the interdependence of deep learning models, we conducted experiments using two key parameters: the power spectrum amplitude ($A_{s}$) and spectral index ($n_{s}$) for CMB temperature power spectra and sky maps. Specifically, we tested 5 values of $A_{s}$ ($2.2\times 10^{-9},~{}2.3\times 10^{-9},~{}2.4\times 10^{-9},~{}2.5\times 10^{-9},~{}2.6\times 10^{-9}$) and 3 values of $n_{s}$ ($0.986,~{}0.996,~{}1.06$). To address the effects of lensing potential on small scales, a high-resolution simulation with $N_{side}=2048$ was utilized. In order to ensure a closer alignment between the simulation results and actual observations, we considered the smoothing effect of the telescope beam on the signal. During the computational phase, we opted for the Gaussian beam as an approximate model for beam shaping to streamline the process.
Figure 4 illustrates the comparison of differences between the CMB T, Q, U sky maps affected by lensing effects and those unaffected by lensing effects. The inputted final dataset for training the U-Net++ model will undergo the influence of the telescope beam effect, which will be manifested as a smoothing effect due to a Gaussian beam with a $\sigma$ value of $0.00311$ rad. Conversely, the corresponding label data will remain unaltered without any modifications.
2.3 The CMB polarization
CMB polarization is measured through time-averaged Stokes parameters – measures of linear polarization of the electric field aligned orthogonally to the line of sight’s (LOS) Cartesian axes. Due to the nature of Thomson scattering, Q and U are ample in describing CMB polarization, given their inability to generate circular polarization (Dodelson & Schmidt, 2021). Astronomical observations reveal that Stoke parameters Q and U are linked by a constant and relative $45^{\circ}$ degree rotation around LOS – whereas the reference frame may revolve freely around it. Bundling the reference frame with a fixed set of normal vectors in rotation will directly link Q and U to the E and B modes.
If we rotate Q and U by an angle $\alpha$ on the plane that is perpendicular to the direction of $\hat{n}$, we obtain the following solution (Zaldarriaga & Seljak, 1997; Zaldarriaga, 1998; Rotti & Huffenberger, 2019),
$(Q\pm iU)^{\prime}(\hat{n})=e^{\mp 2i\alpha}(Q\pm iU)(\hat{n}).$ (11)
We can derive the separation of $E$ and $B$ modes from the Stokes parameters (Kamionkowski et al., 1997a, b; Bunn et al., 2003; Kamionkowski & Kovetz, 2016; Kim & Naselsky, 2010). Here we provide a brief overview of the standard method. Additionally, we can decompose Q and U into ±$\pm 2$ spin spherical harmonics concerning rotation as shown below (Zaldarriaga & Seljak, 1997; Zaldarriaga, 1998),
$Q(\hat{n})\pm iU(\hat{n})=\sum_{\ell,m}a_{\pm 2,\ell m}{{}_{\pm}}{{}_{2}}Y_{\ell m}(\hat{n}),$ (12)
where ${{}_{\pm}}{{}_{2}}Y_{\ell m}(\hat{n})$ are the spin ±$\pm 2$ spherical harmonics, and the coefficients $a_{\pm 2,\ell m}$ are given by
@@ 2.3 The CMB polarization
Based on the theoretical derivation mentioned above, we adopted the same method for simulating temperature sky maps and tested $15$ datasets using two critical parameters, the equation-of-state parameter of dark energy $w$ ($-1.025$, $-1$, $-0.975$) and tensor-to-scalar ratio $r$ (0.001, 0.002, 0.003, 0.004, 0.005, 0.006, 0.007, 0.008, 0.009, 0.01). As a result, we obtained the sky maps of Q-mode and U-mode in real space and showed the difference between the unlensed and lensed sky maps as shown in Figure 4. Similar to the treatment of the temperature field $T$, Gaussian beam smoothing was also applied to Q and U using a $\sigma$ value of $0.00311$. Following this step, we obtained the training dataset for Q and U, with the important caveat that their respective labels remain as unlensed signals.
@@
Figure 2: CMB XX angular power spectrum, $X\in\{E,B\}$. Top panel: The green solid line represents the lensed EE spectrum, while the light green dashed line denotes the unlensed EE spectrum. The blue solid line indicates the lensed BB spectrum, and the light blue dashed line represents the unlensed BB spectrum. The gray solid line depicts the polarization spectrum, encompassing both the EE spectrum and BB spectrum. Bottom panel: The fractional difference between these very similar spectra. The red solid line illustrates the ratio between the unlensed EE spectrum and the lensed EE spectrum, while the pink solid line represents the ratio between the unlensed BB spectrum and the lensed BB spectrum, while the gray solid line represents the reference line where the ratio equals $1$.
Figure 3: The impact of lensing on the CMB TE power spectrum. Top panel: The solid green line depicts the CMB TE angular power spectrum in the absence of lensing. The light green dashed line represents the CMB TE angular power spectrum including the effects of lensing. The gray dashed line illustrates the contribution of noise to the TE angular power spectrum. Bottom panel: The difference between these closely related spectra is shown. The red solid line represents the difference between the TE spectrum without lensing and with lensing, and the gray solid line, as a reference line, represents the position where the difference is equal to 00.
Figure 4: The comparison between the lensed and unlensed CMB maps. The left panel showcases the temperature ($T$) and polarization ($Q,U$) differences on a full sky map with a resolution of $N_{side}=2048$. The map is presented in galactic coordinates, and the units of maps is $\mu K$. On the other hand, the right panel illustrates a patch sky map that focuses on a specific region. This patch map has a size of $512\times 512$, and its center is located approximately at ($281.25,0$) on the left panel. The pixel resolution of the patch map is approximately $1.72$ square arcminutes.
Figure 4: The comparison between the lensed and unlensed CMB maps. The left panel showcases the temperature ($T$) and polarization ($Q,U$) differences on a full sky map with a resolution of $N_{side}=2048$. The map is presented in galactic coordinates, and the units of maps is $\mu K$. On the other hand, the right panel illustrates a patch sky map that focuses on a specific region. This patch map has a size of $512\times 512$, and its center is located approximately at ($281.25,0$) on the left panel. The pixel resolution of the patch map is approximately $1.72$ square arcminutes.
3 U-Net++ Structure and Data Pre-processing
@@
This section introduces the deep neural network architecture U-Net++ (Zhou et al., 2018, 2019) employed in our CMB delensing analysis. In this paper, we utilize the derivative network of U-Net++, called U-Net 2+, as shown in Figure 5. U-Net++ is an enhanced version of the U-Net architecture, incorporating an improved architecture to facilitate the fusion of multiscale features efficiently. The U-Net network is a type of convolutional neural network (CNN) that was initially designed for biomedical image segmentation (Ronneberger et al., 2015). However, it includes considerable structural modifications based on the CNN. The primary objective is to incorporate a sequence of layers to the standard contracting network, with an upsampling operation instead of the pooling operation. As a result, the resolution of the output is increased. The expanded path is relatively symmetrical with the contracted half, generating a U-shaped structure (Ronneberger et al., 2015).
@@
Specifically, the left side of the U-Net is the downsampling (encoder) part, which is used to extract abstract features from the image. By using convolution and downsampling operations, the image size is reduced to extract shallow features. The convolution operation uses valid padding method, ensuring that the results are based on the context features without missing information. Therefore, after each convolution, the size of the image will be reduced.
@@
Figure 5: Evolution from U-Net to U-Net++. Each node in the graph represents a convolution block, downward arrows indicate down-sampling, upward arrows indicate up-sampling, dot arrows indicate skip connections, and the dot box indicates the four outputs. U-Net++ combines U-Nets of different depths into a unified architecture. All substructures share the same encoder, but have their own decoders. Then skip connections are dropped, and every two neighboring nodes are connected with a short skip connection, enabling the deeper decoder to send supervisory signals to the shallower decoder. Finally, by connecting the decoders, a densely connected skip connection is generated so that the dense features propagate along the skip connection, resulting in more flexible feature fusion at the decoder nodes. Thus, each node in the U-Net++ decoder combines multiscale features of the same resolution from all its preceding nodes from a horizontal perspective, and integrates multiscale features of different resolutions from its preceding nodes from a vertical perspective.This multiscale feature aggregation in U-Net++ gradually synthesizes the segmentation, resulting in improved accuracy and fast convergence.
3.1 U-Net++ structure
U-Net++ is a versatile image segmentation architecture developed to enhance image segmentation accuracy. It comprises U-Net++ networks of various depths, with decoders interconnected at the same resolution by means of an improved skip route. By incorporating dense blocks and convolutional layers between the encoder and decoder, it boosts image segmentation accuracy. U-Net++ can be pruned at inference time, if trained with deep supervision, i.e., choose a different output.
@@ 3.1 U-Net++ structure
Figure 5 shows a unified U-Net++ architecture that merges four U-Nets of varying depths. These U-Nets generate four outputs, designated output 1-4. In the graphical abstract, the original U-Net appears as yellow, with skip connections depicted by dot arrows, and the four outputs displayed inside a dotted box. At the inference stage, U-Net++ can be pruned by selecting a different output if it was trained with deep supervision.
U-Net++ employs the same encoder for all its substructures, but each substructure has its unique decoder. The central concept of U-Net++ is to bridge the semantic gap between the feature maps of the encoder and decoder before fusing. To achieve this, a dense convolution block with three convolution layers is used to bridge the semantic gap between ($X^{0,0}$,$X^{1,3}$), for instance.
@@
The proposed technique in this paper aims to accelerate network training by incorporating skip connections. Specifically, by dropping some skip connections and connecting every two neighboring nodes with a short skip connection, the deeper decoder can send supervisory signals to the shallower decoder, leading to faster training. Furthermore, the decoders are connected, creating a densely connected skip connection, which allows dense features to propagate along the skip connection leading to flexible feature fusion. As a result, each node in the U-Net++ decoder combines multi-scale features of the same resolution from all preceding nodes horizontally while integrating multi-scale features of different resolutions from preceding nodes vertically. This multi-scale feature aggregation in U-Net++ gradually synthesizes the segmentation, resulting in improved accuracy and fast convergence.
3.2 Data pre-processing
The U-Net++ model is an evolution of U-Net, sharing many similarities with it. Both models center around semantic segmentation, which detects the class of the object each pixel belongs to for image segmentation (Makinen et al., 2021; Ni et al., 2022). However, it can only process 2D flat images or 3D cubes data, not curved images or sky patches directly.
Figure 6: In the case of $N_{side}=4$, the image presented illustrates a half-orthogonal view projection. In this image, pixels with four neighbors are sequentially labeled in gray, whereas pixels with only three neighbors are highlighted with larger black numbers. It should be noted that only the visible pixels in this view are annotated with their corresponding numbers.
Figure 7: 2D slices from input CMB lensed sky patches (top) and output CMB unlensed sky patches (bottom) for the U-Net++ network. Each patch has a size of $214.86~{}{\rm{deg}}^{2}$. The center of the sky patches is $(l,b)=(101.25^{\circ},19.47^{\circ})$, where $l$ and $b$ are the Galactic longitude and Galactic latitude, respectively. The top row of sky patches is the training dataset, and the bottom row of sky patches is the label. From left to right, the sky patches are T, Q and U map, respectivly. The units of the sky patches is $mK$.
There are currently available network solutions for deep learning on the sphere, such as DeepSphere [^6] https://github.com/deepsphere/deepsphere-cosmo-tf1 (Perraudin et al., 2019) and NNhealpix [^7] https://github.com/ai4cmb/NNhealpix (Krachmalnicoff & Tomasi, 2019). Both methods are highly innovative and have made outstanding contributions to the application of artificial intelligence algorithms on spherical data. We have attempted both methods, but the extremely high resolution of the full sky map we are considering, i.e., $N_{side}=2048$, makes both methods extremely time-consuming during the data initialization stage. Furthermore, we have considered that regardless of the segmentation scheme used, as long as the full sky map is not used as training data, boundary processing issues will arise. Therefore, we require more effective methods for data processing and model training to address this issue. Additionally, due to the fact that not all pixels on HEALPix have exactly four neighboring pixels, but in some cases, there are only three, this may also introduce errors in certain segmentation approaches (Krachmalnicoff & Tomasi, 2019), as shown in Figure 6.
Our segmentation scheme builds upon Makinen’s work (Makinen et al., 2021). Initially, we approximate each pixel at a resolution of $N_{side}=4$) as a flat plane, as shown in Figure 6 with each labeled sky patch. Subsequently, we utilize this as a basis to segment the higher-resolution sky map into patches resembling the $N_{side}=4$ sky map. Ultimately, the high-resolution sky map is divided into 192 flat images (corresponding to the number of pixels $N_{side}=4$).
The U-Net++ architecture requires square image data with HEALPix pixelization, ensuring equal areas for all images. The full-sky map is divided into small sky patches, consisting of 192 equal-area patches with a pixel area of HEALPix $N_{\rm side}$ of 4. Each patch contains $512\times 512$ pixels. In our data structure, a complete sky map is split into (192, 512, 512), as depicted in Figure 7. The sky patches in the top row constitute the training datasets, while the labels are contained in those at the bottom row.
At any observational scale, cosmology inherently encompasses a wealth of physical information. Assuming that we partition an full sky map into multiple independent regions where there is no scale-dependent coupling of information among these regions, any form of segmentation will inevitably entail some degree of informational loss, and our segmentation approach is no exception to this principle. Therefore, we have proposed a mitigation strategy to address this issue.
Firstly, the full sky map is subjected to multiple rotation operations aimed at possible directional dependence issues. Subsequently, each rotated image undergoes segmentation processing. Following this, deep learning techniques are employed to conduct meticulous analysis and reconfiguration of the segmented sky patchs. Thereafter, the processed sky patchs are restored according to their initial rotation angles, and the outcomes obtained from applying the U-Net++ model in multiple iterations are integrated and averaged. Considering our computational resources and time constraints, we have chosen 36 rotation angles. Specifically, $\text{longitude}=[-150^{\circ},-90^{\circ},-30^{\circ},30^{\circ},90^{\circ},150^{\circ}]$ and $\text{latitude}=[-75^{\circ},-45^{\circ},-15^{\circ},15^{\circ},45^{\circ},75^{\circ}]$. This methodology effectively reduces the errors introduced by the segmentation scheme, thereby enhancing the preservation and extraction of a more comprehensive set of cosmological information.
This research proposal aims to mitigate the error impact caused by the segmentation operation to a certain degree, but it cannot completely eliminate it. Additionally, the implementation of multiple image rotation processes necessitates the repetition of the corresponding training process, significantly increasing the demands on computational resources and time costs. Consequently, in our experiments, we limited the number of rotation operations to a total of 35 for the training dataset.
Our method exhibits a considerable degree of adaptability and is well-suited for processing full sky map data utilizing the HEALPix segmentation scheme, such as CMB signals, neutral hydrogen full sky temperature and polarization maps, among others. The specific range of applications encompasses the reconstruction of lensing or CMB signals, the removal of neutral hydrogen foregrounds (Makinen et al., 2021), and the elimination of beam (Ni et al., 2022).
3.3 Loss function

### v3

@@ II.2 The CMB polarization
We can derive the separation of E and B modes from the Stokes parameters Kamionkowski:1996ks; Kamionkowski:1996zd; Bunn:2002df; Kamionkowski:2015yta; kim2010b. Here we provide a brief overview of the standard method. Additionally, we can decompose Q and U into $\pm 2$ spin spherical harmonics concerning rotation as shown below Zaldarriaga:1996xe; zaldarriaga1998cosmic,
@@
Based on the aforementioned theoretical derivations, we employed the same methodology used for simulating temperature sky maps to simulate polarized sky maps. We conducted tests on $30$ datasets utilizing two critical parameters, the equation-of-state parameter of dark energy $w$ (1.025-1.025, 1-1, 0.975-0.975) and tensor-to-scalar ratio $r$ ($0.001$, $0.002$, $0.003$, $0.004$, $0.005$, $0.006$, $0.007$, $0.008$, $0.009$, $0.01$). Similar to the simulation of temperature sky maps, each distinct set of parameters will be employed to generate a corresponding full-sky map.
In the top panel of Figure 1, we use a green solid line and a light green dashed line to show the lensed and unlensed angular power spectra of the E-mode, respectively, and a blue solid line and a light blue dashed line to show the lensed and unlensed angular power spectra of the B-mode, with the equation of state parameter for dark energy w=1w=-1 and the scalar-to-tensor ratio $r=0.005$, respectively. Meanwhile, in the bottom panel, a green dashed line shows the ratio of the lensed to the unlensed E-mode power spectra, and a blue dashed line shows the ratio of the lensed to the unlensed B-mode power spectra. Similar to the simulation of the temperature sky maps, we utilized the Lenspyx tool to generate $30$ sky maps corresponding to the polarization power spectra.
After obtaining the full-sky maps of the Stokes parameters T, Q, and U, we transformed them into the T, E, and B components in harmonic space using Equation (12). This conversion enables a more direct analysis of the physical polarization modes and separates the scalar (E-mode) and tensor (B-mode) contributions. The resulting T/E/B maps serve as the input for our neural network, which is specifically designed to operate on the T/E/B fields for subsequent data processing and training.
@@
Figure 1: CMB $X$ angular power spectrum, $X\in\{{\rm T},{\rm E},{\rm B}\}$. Top Panel: The red solid line represents the lensed TT spectrum, while the light red dashed line corresponds to the unlensed TT spectrum. The green solid line denotes the EE spectrum under the influence of lensing, whereas the light green dashed line shows the EE spectrum without lensing effects. The blue solid line indicates the lensing-modified BB spectrum, and the light blue dashed line depicts the BB spectrum in the absence of lensing effects. The parameters for generating the temperature angular power spectrum are $A_{s}=2.1\times 10^{-9}$ and $r=0.005$. The black solid line displays the noise level of the temperature detectors, and the grey solid line represents that of the polarization detectors. Bottom Panel: It illustrates the relative differences between these highly similar spectra. The red dashed line marks the ratio of the unlensed TT spectrum to the lensed TT spectrum; the green dashed line indicates the ratio of the unlensed $E$ spectrum to the lensed $E$ spectrum; and the blue dashed line shows the ratio of the unlensed BB spectrum to the lensed $B$ spectrum.
II.3 Noise simulation
The noise in CMB detectors is often approximated as Gaussian white noise, and its angular power spectral density can be defined in the following Wu:2014hta; BICEP2:2015nss; wu2020detecting; Wolz:2023lzb:
${\rm N}_{\ell}^{XX^{\prime}}=[s^{2}e^{(-\ell(\ell+1)\frac{\theta_{{\rm FWMH}}^{2}}{8\ln 2})}]^{-1},$ (14)
where $s$ denotes the telescope’s sensitivity, $X\in\{{\rm T},{\rm E},{\rm B}\}$, $\theta_{{\rm FWHM}}$ represents the FWHM of the telescope. For simplicity, we adopt $\theta_{{\rm FWHM}}=7.9~{\rm arcmin}$, with a sensitivity for the temperature detector of $1.5~{\rm\mu K}$ arcmins and for the polarization detector of $2.1~{\rm\mu K}$ arcmins NASAPICO:2019thw.
According to Equation (14), we can calculate the temperature and polarization angular power spectra of noise, as illustrated by the black and grey lines in the top panel of Figure 1. For the noise of each component, we generate corresponding full-sky maps based on their angular power spectra and overlay these noise maps onto the sky maps with lensing effects. Similar to the simulation of the temperature sky maps, we utilized the Lenspyx tool to generate $30$ sky maps corresponding to the noise power spectra.
II.4 Data pre-processing
By conducting simulations with Lenspyx, we generated 30 sets of full-sky maps for I, Q, U, along with noise maps. Using Formula 1, we were then able to obtain the sky maps corresponding to our training data, i.e. T, E, B fields. Spherical data poses challenges for deep learning due to its non-Euclidean structure. Traditional convolutions are not directly applicable on the sphere because of the lack of translational equivariance, and the rotational symmetry on the sphere is also difficult to handle. Moreover, spherical sampling is often irregular, such as in the HEALPix grid, which complicates feature extraction.
Figure 2: A half-orthogonal view projection of the sky map.Gray sequentially labels pixels with four neighbors, while larger black numbers highlight those with only three.
Figure 3: The T, E and B sky map patch. Each patch has a size of $214.86~{{\rm deg}}^{2}$. The center of the sky patches is $(l,b)=(101.25{\rm{}^{\circ}},19.471{\rm{}^{\circ}})$, where $l$ and $b$ are the Galactic longitude and Galactic latitude, respectively. From left to right, the sky patches are T, E and B map, respectivly. The top row of patches displays sky map patches affected by lensing effects, including noise and instrumental artifacts, which are used to construct the training dataset. The second row presents the original, unlensed sky map patches, which serve as the basis for generating the label dataset. The unit is $\rm\mu K$}.
There are currently available network solutions for deep learning on the sphere, such as DeepSphere Perraudin:2018rbt and NNhealpix Krachmalnicoff:2019zjh. Both methods are highly innovative and have made outstanding contributions to the application of artificial intelligence algorithms on spherical data. We have attempted both methods, but the extremely high resolution of the full sky map we are considering, i.e., $N_{\rm side}=2048$, makes both methods extremely time-consuming during the data initialization stage. However, regardless of the segmentation scheme used, if the full-sky map is not directly used as training data, boundary processing issues will inevitably arise. Therefore, we require more effective data processing and model training methods to address this challenge. Additionally, due to the fact that not all pixels on HEALPix have exactly four neighboring pixels, but in some cases, there are only three, this may also introduce errors in certain segmentation approaches Krachmalnicoff:2019zjh, as shown in Figure 2.
When there is a mismatch between data and model, there are usually two ways to deal with it: modifying the model to fit the data, or adjusting the data format to meet the model requirements. In related research such as DeepSphere and NNhealpix, the first strategy is used, which is to modify the structure of the deep learning model to adapt to spherical data, so that the network can complete training. However, there are certain limitations in such methods, so we propose an alternative approach that segments spherical data to more flexibly address the mismatch between data and models. Our segmentation scheme builds upon Makinen’s work Makinen:2020gvh. Specifically, we approximate each pixel at a resolution of $N_{\rm side}=4$ as a flat region, as illustrated by the labeled sky patches in Figure 2. Consequently, the entire sky is divided into 192 independent small sky region images.
For simulated sky maps with a resolution of $N_{\rm side}=2048$, after applying the aforementioned segmentation method, each full sky map is broken down into $192$ image patches (number of pixels at resolution $N_{\rm side}=4$), each sized $512\times 512$ pixels, as shown in Figure 3. The sky patches in the first row are used to form the training dataset, while the corresponding labels are contained within the sky patches in the second row.
At any observational scale, cosmology inherently encompasses a wealth of physical information. Assuming that we partition an full sky map into multiple independent regions where there is no scale-dependent coupling of information among these regions, any form of segmentation will inevitably entail some degree of informational loss, and our segmentation approach is no exception. Therefore, we have proposed a mitigation strategy to address this issue.
We constructed a technical framework based on a pixel-level averaging fusion method, consisting of the following key steps: First, the original full-sky CMB map is randomly rotated by predefined angles. The rotated full-sky map is then divided into map patches. Subsequently, an convolutional neural network (CNN) network architecture is applied to perform delensing on each segmented image patch, generating corresponding intermediate delensed results. After the forward processing is completed, inverse rotation operations are performed on each intermediate result to restore the original spatial orientation.
By repeating this entire process 30 times, and given that the segmentation edges vary in each iteration due to the specific rotations, the final aggregated result effectively alleviates the errors introduced by any single segmentation, thereby significantly mitigating the errors caused by the segmentation process.
This operation is applied independently to each field of the sky maps, requiring 30 rotational iterations per sky map type with corresponding repetition of the training process. The rotation angles are configured with $\phi$ values are $180{\rm{}^{\circ}},-120{\rm{}^{\circ}},-60{\rm{}^{\circ}},0{\rm{}^{\circ}},60{\rm{}^{\circ}},120{\rm{}^{\circ}}]$, and $\theta$ values are $60{\rm{}^{\circ}},-30{\rm{}^{\circ}},-0{\rm{}^{\circ}},30{\rm{}^{\circ}},60{\rm{}^{\circ}}]$.
The rotation scheme aims to mitigate the error impact caused by the segmentation operation to a certain degree, but it cannot completely eliminate it. Additionally, the implementation of multiple image rotation processes necessitates the repetition of the corresponding training process, significantly increasing the demands on computational resources and time costs. Consequently, in our research, we limited the number of rotation operations to a total of $30$ for the training dataset.
It is worth noting that since E-mode and B-mode represent curl-free and divergence-free polarization fields respectively, they can effectively separate signals of different physical origins. In particular, B-mode produced by primordial gravitational waves and those induced by gravitational lensing effects can be more clearly distinguished and processed in T/E/B space. Therefore, using the T/E/B representation for CMB delensing is the optimal choice. After obtaining the I, Q, U Stokes parameters that have been affected by noise and instrumental effects, we can derive the T, E, and B mode sky maps using Equation (4) and Equation (12).
III QE Delensing and UNet++ Structure
In the field of cosmic signal analysis and intelligent image processing, the quadratic estimator (QE) algorithm, a secondary estimator, relies on mathematical statistics to mine deep correlations in observed signals, aiding in the inversion of cosmological parameters Okamoto:2003zw; Schaan:2018tup; Belkner:2023duz. The UNet++ algorithm achieves high-precision image segmentation through deep convolutional networks and multi-scale feature fusion zhou2018unet++; zhou2019unet++. Although the two have different paths, the former is based on physical statistical modeling, while the latter relies on data-driven learning, they jointly build a core methodology for extracting key information from complex data. We subsequently applied these two methods to delens CMB and then compared the discrepancies in their processing outcomes.
III.1 QE delensing
The QE delensing is a powerful technique that reconstructs and removes the gravitational‐lensing signal imprinted on the CMB by large‐scale structure, thereby sharpening acoustic features and reducing spurious B-mode power. At its core, this method uses optimal quadratic combinations of temperature (T) and polarization (E, B) multipoles to estimate the lensing potential, then “undoes” the inferred deflections to recover a closer‐to‐primordial map Diego-Palazuelos:2020lme.
In practice, the QE algorithm reconstructs the lensing potential $\phi$ by exploiting the off-diagonal covariance induced by gravitational deflection. This deflection introduces correlations between originally independent Fourier modes of the unlensed CMB Hu:2001kj. Schematically, one forms weighted pairs of observed multipoles and integrates over $\ell$ to yield Bohm:2016gzt
$\hat{\phi}_{XY}(L)=N_{L}^{XY}\int\frac{{\rm d}^{2}l}{(2\pi)^{2}}X(l)Y(L-l)f(l,L)$ (15)
where $X,Y\in{{\rm T},{\rm E},{\rm B}}$ and the filter $f^{XY}$ depends on theoretical $C_{\ell}^{\rm XY}$. The deflection field ${\rm\nabla}\phi$ mixes the unlensed E- and B-polarizations.
The normalization $N_{L}$ is chosen so that $\langle\phi^{{\rm TT}}\rangle=\phi$. In practice one combines all channel pairs into a minimum-variance $\hat{\phi}(L)$. Analogous filters $f^{XY}$ are derived for all pairs; a weighted sum of them yields the minimum-variance estimate $\hat{\phi}(L)$ Hu:2001kj.
We will perform delensing operations on all-sky temperature and polarization maps using the delensalot package, and conduct a systematic quantitative analysis to compare the delensing efficacy with that of the UNet++ architecture Belkner:2023duz.
III.2 UNet++ structure
@@
UNet++ is an enhanced version of the UNet architecture, incorporating an improved architecture to facilitate the fusion of multiscale features efficiently. The UNet network is a type of CNN that was initially designed for biomedical image segmentation ronneberger2015u. However, it includes considerable structural modifications based on the CNN. The primary objective is to incorporate a sequence of layers to the standard contracting network, with an upsampling operation instead of the pooling operation. As a result, the resolution of the output is increased. The expanded path is relatively symmetrical with the contracted half, generating a U-shaped structure ronneberger2015u.This section introduces the deep neural network architecture UNet++ zhou2018unet++; zhou2019unet++ employed in our CMB delensing analysis. In this paper, we utilize the derivative network of UNet++, as shown in Figure 4. UNet++ is an enhanced version of the UNet architecture, incorporating an improved architecture to facilitate the fusion of multiscale features efficiently. The UNet network is a type of CNN that was initially designed for biomedical image segmentation ronneberger2015u. However, it includes considerable structural modifications based on the CNN. The primary objective is to incorporate a sequence of layers to the standard contracting network, with an upsampling operation instead of the pooling operation. As a result, the resolution of the output is increased. The expanded path is relatively symmetrical with the contracted half, generating a U-shaped structure ronneberger2015u.
@@
Specifically, the left side of the UNet is the downsampling (encoder) part, which is used to extract abstract features from the image. By using convolution and downsampling operations, the image size is reduced to extract shallow features. The convolution operation uses valid padding method, ensuring that the results are based on the context features without missing information. Therefore, after each convolution, the size of the image will be reduced.
@@
Figure 4: Unet++ network architecture. Each node in the graph represents a convolution block, downward arrows indicate down-sampling, upward arrows indicate up-sampling, dot arrows indicate skip connections, and the dot box indicates the four outputs. UNet++ combines UNets of different depths into a unified architecture. All substructures share the same encoder, but have their own decoders. Then skip connections are dropped, and every two neighboring nodes are connected with a short skip connection, enabling the deeper decoder to send supervisory signals to the shallower decoder. Finally, by connecting the decoders, a densely connected skip connection is generated so that the dense features propagate along the skip connection, resulting in more flexible feature fusion at the decoder nodes. Thus, each node in the UNet++ decoder combines multiscale features of the same resolution from all its preceding nodes from a horizontal perspective, and integrates multiscale features of different resolutions from its preceding nodes from a vertical perspective.This multiscale feature aggregation in UNet++ gradually synthesizes the segmentation, resulting in improved accuracy and fast convergence.
@@
Figure 4 shows a unified UNet++ architecture that merges four UNets of varying depths. These UNets generate four outputs, designated output 1-4. In the graphical abstract, the original UNet appears as yellow, with skip connections depicted by dot arrows, and the four outputs displayed inside a dotted box. At the inference stage, UNet++ can be pruned by selecting a different output if it was trained with deep supervision.
@@
By dropping some skip connections and connecting every two neighboring nodes with a short skip connection, the deeper decoder can send supervisory signals to the shallower decoder, leading to faster training. Furthermore, the decoders are connected, creating a densely connected skip connection, which allows dense features to propagate along the skip connection leading to flexible feature fusion. As a result, each node in the UNet++ decoder combines multi-scale features of the same resolution from all preceding nodes horizontally while integrating multi-scale features of different resolutions from preceding nodes vertically. This multi-scale feature aggregation in UNet++ gradually synthesizes the segmentation, resulting in improved accuracy and fast convergence.
III.3 Loss function